OpenStudy (ttop0816):

Simplify. csc^2θ + cot^2θ - 1 0 2 cot^2θ 2cot^2θ

OpenStudy (anonymous):

\[ \cot ^2(\theta )+\csc ^2(\theta )-1= cot ^2(\theta )+ cot ^2(\theta )+1 -1\\ \]

OpenStudy (anonymous):

It is the last choice

OpenStudy (ybarrap):

$$ \cot^2\theta= \csc^2θ -1\\ $$ So $$ \csc^2θ + \cot^2θ - 1\\ =\cot^2\theta+cot^2\theta\\ =2\cot^2\theta $$